# Supplementary Materials

### v2

@@ I.1 The Model
Assumption 2: Initial researcher productivity $q_{0}$ follows an exponential distribution (LABEL:fig:criteriaA).
Assumption 3: The change in productivity from year to year $\delta_{t}=q_{t+1}-q_{t}$ is distributed independently of previous years’ productivity, and follows a Laplace distribution (LABEL:fig:criteriaB-D).
Assumption 4: Researcher productivity exhibits different dynamics in different career stages (LABEL:fig:criteriaE).
@@
Simplified model. We fix the mode of the Laplace distribution globally to an arbitrary constant ($\mu=-1$) and fit the first-year exponential rate parameter using the empirical mean of the data, $\widehat{\lambda}_{0}=4.65$. Then we separate the data into two career stages: years 00 to $4$, representing the “early career” stage, and years $5$ to $20$, representing the remainder of a career. Each career stage has the same location $\mu=-1$, but different scale parameters $\alpha_{1}$ and $\alpha_{2}$, respectively, which are the only free parameters of this simplified model. We then simulate trajectories from the simplified model, and assess whether individual simulated trajectories exhibits the canonical pattern using a simple model selection approach across a family of linear and piece-wise linear regressions. We systematically explore the variance parameter space of this model, and then plot the fraction of trajectories that meet the criteria for being canonical to produce LABEL:fig:simplified_canonical.
We follow Ref. [way2017misleading] in using model selection to classify individual trajectories as canonical. We fit a two-part piecewise linear model to each trajectory for each possible change point between career years 3 and 17, as well as a linear model. We then use the small-sample correction of the AIC (AICc) to determine the best-fitting model. The trajectory is labeled as canonical if the best-fitting model is a piecewise model (that is, the linear model is not the best fit) that additionally fulfills the following criteria: the slope of the first piece is positive, the slope of the second piece is negative, and the magnitude of the slope of the first piece is at least twice the magnitude of the slope of the second piece.
@@
A career stage $i$ is defined in terms of change points $c_{i},c_{i+1}$, where we construct the data within each career stage $D_{i}=\{(q_{t},q_{t+1})\}$, where $c_{i}\leq t<c_{i+1}$. Each candidate set of career stages is identified by a tuple of change points $(c_{1},c_{2},c_{3})$, where $c_{3}$ or both $c_{2}$ and $c_{3}$ may be omitted. We implicitly assume a change point at the two endpoints for career age 00 and at $t_{\text{max}}$, and we adopt the convention that the career stages include the right endpoints. For instance, the change point set $(2,5)$ would encode career stages 0–2, 3–5, and 6–20. This procedure yields 1,159 possible sets of one, two, or three change points from 1 to 19.
@@
We estimate a single exponential parameter $\widehat{\lambda}_{0}>0$ for the distribution of $q_{0}$, which is shared across all models and is independent of the choice of change points. To generate a synthetic productivity trajectory, we first generate a choice of initial productivity $q_{0}$ from the estimated model $\mathbb{P}(q_{0}|\lambda_{0})$. Then we set $q_{t+1}=q_{t}+\delta_{t}$ for $c_{i}<t\leq c_{i+1}$ which we draw from the parameterized distribution $\mathbb{P}(\delta_{t}|q_{t},\alpha_{i},\beta_{i})$. The synthetic productivity trajectories simulated from the model are used to perform the comparisons with the empirical data shown in LABEL:fig:fit and LABEL:fig:unfit.
I.2 Validity of the Markovian assumption
@@
$P(q_{t+1}=r_{t+1}$ $|q_{t}=r_{t},q_{t-1}=r_{t-1})$ (1)
$=P(q_{t+1}=r_{t+1}$ $|q_{t}=r_{t}).$ (2)
By conditioning on the intermediate value $q_{t}$, we can observe whether $\mathbb{P}(q_{t+1})$ varies by the value in the previous timestep, $q_{t-1}$ (Fig. S1). We find confirmation for two of the results stated in the main paper. First, compared to simulated trajectories, the empirical trajectories have less variance within careers, greater variance across careers, and many more years with zeros (LABEL:fig:unfit). Here we confirm this observation from the point of view of the Markov property: the overall trend of increasing means implies that historically high (resp. low) values lead future values to be high (resp. low), even after conditioning on intermediate timesteps (Fig. S1). In the main paper, we find significantly greater abundance of years with zero papers empirically than our model predicted, even after correcting for missingness in the data (LABEL:fig:unfitB,D). We still see that the non-Markovian behavior is strongest around zero (that is, conditioning on $q_{t}=0$, we see that trajectories where $q_{t-1}=0$ are far more likely to produce zeros in year t+1 than other values of $q_{t-1}$), but nevertheless there is a smaller non-Markovian trend in the other conditional distributions as well, particularly in the top-most quantiles. This provides validation for our explanation for how empirical trajectories reach higher cumulative productivities by year 5 than the simulations would predict (LABEL:fig:unfitC).
@@
I.3 The distribution of the random walk increments
@@
II Recovering known parameters from synthetic data
@@
Using synthetic trajectories with known parameters, the full model correctly recovers the relevant parameters generally up to an error of ±$\pm 0.1$ for $\lambda$ and $\alpha$, and ±$\pm 0.01$ for $\beta$ (Fig. S3). To put that in perspective, an error in $\lambda$ of $0.1$ corresponds to a tenth of a publication in a scientist’s first year, and the errors in $\alpha$ and $\beta$ can be interpreted on a similar scale. This scale of error is practically negligible compared to the much larger underlying variability in productivity across individuals.
@@ III DBLP Data
This random walk model can be applied to any dataset composed of time series of individual researcher productivities that range onward from the beginning of a career, e.g., the first year of a permanent research position like a tenure-track job at a PhD-granting institution. Datasets of individual-level productivity derived from bibliographic databases like the Web of Science [clarivate2022wos] are attractive because of their scale, but pose additional complexities due to the need to first disambiguate individuals, then stratify by fields, which can exhibit widely different average productivity levels [piro2013macro], and finally stratify by different roles, e.g., faculty vs. trainees, or researchers employed at institutions with different research intensities.
@@
We avoid such complexities by studying a dataset of known computer science faculty at PhD-granting institutions in the US or Canada [way2017misleading], which we linked to their publications as recorded in DBLP [dblp2016november], a bibliographic database that focuses on computing and which was used to identify the underlying diversity of productivity trajectories [way2017misleading]. The faculty dataset was a complete collection of the 205 department or school-level academic units on the Computing Research Association’s Forsythe List of PhD-granting departments in computing-related disciplines in the US and Canada [clauset2015systematic, cra2012forsythe]. The manual collection process yielded 5,032 faculty in these units, and the dataset was subsetted to the 2,583 whose PhD degree and first assistant professorship appointment were at units in the sample. The DBLP was then joined to the faculty census, using manual name disambiguation as necessary, yielding 2,453 faculty with linked publication records. The further inclusion criteria described in the “Data” section of the main manuscript resulted in the reported 2,085 faculty in our analysis.
@@
Current faculty often begin publishing before the start of their faculty career, but the underlying dynamics of productivity are different in the pre-faculty stage of a career [zhang2022labor]. We focus our analysis here only on productivity during a faculty career, and so we exclude those publications from years prior to that career’s beginning. Although DBLP has reasonably good coverage over most venues in which computer science faculty publish, it is not complete, and faculty working in subfields that are not well-indexed by DBLP, such as interdisciplinary computing, would appear as anomalously unproductive researchers.
@@
In order to correct for rising individual productivity and uneven historical coverage, Way et al. adjusted historical productivities using a linear model estimated from publication data manually extracted from CVs [way2017misleading]. Such a linear adjustment preserves years with zero papers, which may lead to an over-estimate in the number of years with zero publications. To evaluate this possibility, we randomly select eight individuals among those with at least ten years of zero publications, and manually counted the number of years when they have zero publications in both DBLP (108) and on their websites and Google Scholar (82). This tally yields a maximum likelihood binomial estimate that $75.9\%$ (95% CI: $(67.9\%,84.0\%)$) of DBLP-observed zero-publication years are correct. With the most conservative correction ($67.9\%$ of years with zero publications are truly zero), and most conservative estimate for the empirical mean number of zeros ($14.1\%$ at the bottom of the estimated 95% CI), we would still find that $9.5\%$ of years in the full empirical trajectories are years with zero publications. This estimate excludes the 95% CI of the mean number of zeros in 10,000 simulated trajectories $(8.5\%,8.7\%)$, implying that zero-productivity years are likely more common in real productivity trajectories than our model can account for.

### v3

# Scientific productivity as a random walk

@@ IX.1 The Model
Assumption 2: Initial researcher productivity $q_{0}$ follows an exponential distribution (Fig. 1A).
Assumption 3: The change in productivity from year to year $\delta_{t}=q_{t+1}-q_{t}$ is distributed independently of previous years’ productivity, and follows a Laplace distribution (Fig. 1B-D).
Assumption 4: Researcher productivity exhibits different dynamics in different career stages (Fig. 1E).
@@
Simplified model. We fix the mode of the Laplace distribution globally to an arbitrary constant ($\mu=-1$) and fit the first-year exponential rate parameter using the empirical mean of the data, $\widehat{\lambda}_{0}=4.65$. Then we separate the data into two career stages: years 0 to $4$, representing the “early career” stage, and years $5$ to $20$, representing the remainder of a career. Each career stage has the same location $\mu=-1$, but different scale parameters $\alpha_{1}$ and $\alpha_{2}$, respectively, which are the only free parameters of this simplified model. We then simulate trajectories from the simplified model, and assess whether individual simulated trajectories exhibits the canonical pattern using a simple model selection approach across a family of linear and piece-wise linear regressions. We systematically explore the variance parameter space of this model, and then plot the fraction of trajectories that meet the criteria for being canonical to produce Fig. 2.
We follow Ref. [57] in using model selection to classify individual trajectories as canonical. We fit a two-part piecewise linear model to each trajectory for each possible change point between career years 3 and 17, as well as a linear model. We then use the small-sample correction of the AIC (AICc) to determine the best-fitting model. The trajectory is labeled as canonical if the best-fitting model is a piecewise model (that is, the linear model is not the best fit) that additionally fulfills the following criteria: the slope of the first piece is positive, the slope of the second piece is negative, and the magnitude of the slope of the first piece is at least twice the magnitude of the slope of the second piece.
@@
A career stage $i$ is defined in terms of change points $c_{i},c_{i+1}$, where we construct the data within each career stage $D_{i}=\{(q_{t},q_{t+1})\}$, where $c_{i}\leq t<c_{i+1}$. Each candidate set of career stages is identified by a tuple of change points $(c_{1},c_{2},c_{3})$, where $c_{3}$ or both $c_{2}$ and $c_{3}$ may be omitted. We implicitly assume a change point at the two endpoints for career age 0 and at $t_{\text{max}}$, and we adopt the convention that the career stages include the right endpoints. For instance, the change point set $(2,5)$ would encode career stages 0–2, 3–5, and 6–20. This procedure yields 1,159 possible sets of one, two, or three change points from 1 to 19.
@@
We estimate a single exponential parameter $\widehat{\lambda}_{0}>0$ for the distribution of $q_{0}$, which is shared across all models and is independent of the choice of change points. To generate a synthetic productivity trajectory, we first generate a choice of initial productivity $q_{0}$ from the estimated model $\mathbb{P}(q_{0}|\lambda_{0})$. Then we set $q_{t+1}=q_{t}+\delta_{t}$ for $c_{i}<t\leq c_{i+1}$ which we draw from the parameterized distribution $\mathbb{P}(\delta_{t}|q_{t},\alpha_{i},\beta_{i})$. The synthetic productivity trajectories simulated from the model are used to perform the comparisons with the empirical data shown in Fig. 3 and Fig. 4.
IX.2 Validity of the Markovian assumption
@@
$P(q_{t+1}=r_{t+1}$ $|q_{t}=r_{t},q_{t-1}=r_{t-1})$ (S1)
$=P(q_{t+1}=r_{t+1}$ $|q_{t}=r_{t}).$ (S2)
By conditioning on the intermediate value $q_{t}$, we can observe whether $\mathbb{P}(q_{t+1})$ varies by the value in the previous timestep, $q_{t-1}$ (Fig. S3). We find confirmation for two of the results stated in the main paper. First, compared to simulated trajectories, the empirical trajectories have less variance within careers, greater variance across careers, and many more years with zeros (Fig. 4). Here we confirm this observation from the point of view of the Markov property: the overall trend of increasing means implies that historically high (resp. low) values lead future values to be high (resp. low), even after conditioning on intermediate timesteps (Fig. S3). In the main paper, we find significantly greater abundance of years with zero papers empirically than our model predicted, even after correcting for missingness in the data (Fig. 4B,D). We still see that the non-Markovian behavior is strongest around zero (that is, conditioning on $q_{t}=0$, we see that trajectories where $q_{t-1}=0$ are far more likely to produce zeros in year t+1 than other values of $q_{t-1}$), but nevertheless there is a smaller non-Markovian trend in the other conditional distributions as well, particularly in the top-most quantiles. This provides validation for our explanation for how empirical trajectories reach higher cumulative productivities by year 5 than the simulations would predict (Fig. 4C).
@@
IX.3 The distribution of the random walk increments
@@
X Recovering known parameters from synthetic data
@@
Using synthetic trajectories with known parameters, the full model correctly recovers the relevant parameters generally up to an error of $\pm 0.1$ for $\lambda$ and $\alpha$, and $\pm 0.01$ for $\beta$ (Fig. S5). To put that in perspective, an error in $\lambda$ of $0.1$ corresponds to a tenth of a publication in a scientist’s first year, and the errors in $\alpha$ and $\beta$ can be interpreted on a similar scale. This scale of error is practically negligible compared to the much larger underlying variability in productivity across individuals.
@@ XI DBLP Data
This random walk model can be applied to any dataset composed of time series of individual researcher productivities that range onward from the beginning of a career, e.g., the first year of a permanent research position like a tenure-track job at a PhD-granting institution. Datasets of individual-level productivity derived from bibliographic databases like the Web of Science [8] are attractive because of their scale, but pose additional complexities due to the need to first disambiguate individuals, then stratify by fields, which can exhibit widely different average productivity levels [45], and finally stratify by different roles, e.g., faculty vs. trainees, or researchers employed at institutions with different research intensities.
@@
We avoid such complexities by studying a dataset of known computer science faculty at PhD-granting institutions in the US or Canada [57], which we linked to their publications as recorded in DBLP [15], a bibliographic database that focuses on computing and which was used to identify the underlying diversity of productivity trajectories [57]. The faculty dataset was a complete collection of the 205 department or school-level academic units on the Computing Research Association’s Forsythe List of PhD-granting departments in computing-related disciplines in the US and Canada [9, 1]. The manual collection process yielded 5,032 faculty in these units, and the dataset was subsetted to the 2,583 whose PhD degree and first assistant professorship appointment were at units in the sample. The DBLP was then joined to the faculty census, using manual name disambiguation as necessary, yielding 2,453 faculty with linked publication records. The further inclusion criteria described in the “Data” section of the main manuscript resulted in the reported 2,085 faculty in our analysis.
@@
Current faculty often begin publishing before the start of their faculty career, but the underlying dynamics of productivity are different in the pre-faculty stage of a career [59]. We focus our analysis here only on productivity during a faculty career, and so we exclude those publications from years prior to that career’s beginning. Although DBLP has reasonably good coverage over most venues in which computer science faculty publish, it is not complete, and faculty working in subfields that are not well-indexed by DBLP, such as interdisciplinary computing, would appear as anomalously unproductive researchers.
@@
In order to correct for rising individual productivity and uneven historical coverage, Way et al. adjusted historical productivities using a linear model estimated from publication data manually extracted from CVs [57]. Such a linear adjustment preserves years with zero papers, which may lead to an over-estimate in the number of years with zero publications. To evaluate this possibility, we randomly select eight individuals among those with at least ten years of zero publications, and manually counted the number of years when they have zero publications in both DBLP (108) and on their websites and Google Scholar (82). This tally yields a maximum likelihood binomial estimate that $75.9\%$ (95% CI: $(67.9\%,84.0\%)$) of DBLP-observed zero-publication years are correct. With the most conservative correction ($67.9\%$ of years with zero publications are truly zero), and most conservative estimate for the empirical mean number of zeros ($14.1\%$ at the bottom of the estimated 95% CI), we would still find that $9.5\%$ of years in the full empirical trajectories are years with zero publications. This estimate excludes the 95% CI of the mean number of zeros in 10,000 simulated trajectories $(8.5\%,8.7\%)$, implying that zero-productivity years are likely more common in real productivity trajectories than our model can account for.
XII Skewness and kurtosis of productivity increments
To further diagnose differences between the empirical DBLP data and the simulated full model output, we computed the skewness and kurtosis within each of the three inferred career stages for both empirical and simulated trajectories, constructing 95% block bootstrap CIs at the person level for the DBLP data, and regular 95% bootstrap CIs for the simulated trajectories. We found skewness to be consistently higher in simulated than empirical trajectories. In the first career stage from year 1 to 4, the skewness was $0.232$ (95% CI [$0.133$, $0.327$]) for empirical data and $0.765$ (95% CI [$0.696$, $0.834$]) for simulated data; from years 5–7, the skewness was $0.141$ (95% CI [$0.024$, $0.246$]) empirically and $0.689$ (95% CI [$0.606$, $0.776$]) for simulation; and from years 8–end, the skewness was $0.034$ (95% CI [0.053-0.053, $0.123$]) empirically and $0.721$ (95% CI [$0.679$, $0.760$]) for simulation. By contrast, kurtosis did not exhibit a consistent pattern: it was greater in the model than in the data during the first two stages, but the reverse held in the final stage. Specifically, in years 1–4, empirical kurtosis was $2.292$ (95% CI [$1.827$, $2.662$]) compared to $2.998$ (95% CI [$2.670$, $3.332$]) in the model; in years 5–7, it was $2.360$ (95% CI [$1.735$, $2.834$]) empirically versus $2.941$ (95% CI [$2.512$, $3.507$]) in the model; and in years 8–end, it was $4.024$ (95% CI [$3.302$, $4.599$]) empirically compared to $2.929$ (95% CI [$2.724$, $3.143$]) in the model.
XIII Connection to Shockley model of productivity
Our model results in the surprising recapitulation of a pattern on research productivity introduced by William Shockley in 1957 [49]. Shockley found that over the course of a variety of scientific careers, the cumulative productivity within institutions tended toward a lognormal distribution, and had theorized some possible mechanisms, namely, that scientific papers are the result of combining individual traits or skills in such a way that the success of the project depends on each of these traits or skills. In other words, these traits or skills are connected in series, so that the outcome depends multiplicatively on each of these components. Then, if these traits or skills were each normally distributed, their product would be lognormally distributed.
Figure S6: Cumulative productivity distributions compared to fitted normal distributions. (A) Cumulative productivities for 10,000 simulated careers of 21 years from full random walk models compared with fitted normal distribution, and 2-sided KS test statistic with p-value, using the dataset of all DBLP trajectories. (B) Cumulative productivities and fitted normal distributions of all empirical trajectories.
Interestingly, productivity trajectories simulated from our full model also produce cumulative productivities over a 21-year span that are statistically indistinguishable from a lognormal distribution (Fig. S6A). By contrast, the empirical data is not immediately lognormal, due to a surfeit of less productive careers (Fig. S6B). The same pattern holds when using simulations from the model only fitted to full 21-year-long careers (the full data; Fig. S7). Shockley also found a wider tail than lognormal empirically, however, which he overcame through manually selecting a cut-off through “trial and error”.
Our model poses potential theoretical challenges to Shockley’s model. Shockley anticipated and argued against the “organizational hypothesis” that coauthorship produces the lognormal distributions of cumulative productivity. However, he did not consider the possibility that identically skilled individuals can achieve cumulative productivities that are lognormally distributed simply through the variability accrued by careers progressing through multiple career stages. Indeed, our model presents a potential alternative to Shockley’s preferred explanation of differences in individual aptitude.
Figure S7: Cumulative productivity distributions compared to fitted normal distributions. (A) Cumulative productivities for simulated careers from full random walk models compared with fitted normal distribution, and 2-sided KS test statistic with p-value, using the dataset of all DBLP trajectories. (B) Cumulative productivities for simulated careers from full random walk models compared with fitted normal distribution, and 2-sided KS test statistic with p-value, only using the dataset of full DBLP trajectories.
XIV Baseline models
XIV.1 Exponential distribution of papers around time-varying mean
Here we motivate our random walk approach model by directly mathematizing a model posited in the “canonical trajectory” literature: the “aging functions” introduced by Ref. [3]. In particular, we model a time-varying mean with exponential distribution of papers around that mean, with independent productivity across timesteps conditioned on the mean. We note that the difference between two iid exponential distributions is a Laplace distribution, providing the model with some initial credibility.
We model to changing mean following Ref. [57], by assuming that the average trajectory is either linear or piecewise-linear with two components. We choose between the linear model or the specific piecewise-linear model using model selection. In particular, we fit a piecewise linear model with two pieces on every possible changepoint between years 3 and 17 for the entire dataset of full 21-year trajectories, as well as a linear model without any changepoint, and select the model out of that family with the best AICc.
Figure S8: Minimal model with exponentially-distributed productivity around a time-varying mean. Minimal model with exponentially-distributed productivity around a time-varying mean. (A) Average career productivity for each career age, simulated from the fitted model. The vertical dotted line indicates the inferred changepoint, and error bands indicate 95% CIs. (B) Distributions of years with greatest productivity within each (blue) empirical and (orange) simulated trajectory. Histograms are normalized such that each bar represents the probability that a given trajectory has its year of greatest productivity within that bin.
The best fitting model displayed a changepoint at the fourth career year, with initial intercept at $4.77$ papers and pre-year-4 slope of $0.98$ more papers/year (Fig. S8A). After year 4, the slope changes to 0.06-0.06. As expected, this fulfills the criteria from Ref. [57] of being a canonical trajectory, since the first slope is positive, the second slope is negative, and the absolute value of the first slope is at least twice that of the absolute value of the second slope. We ignore the fact that the simulated average trajectory does not look precisely like the empirical data, since through semi-parametric approaches, one could theoretically fit the average trajectory as closely as one desires. More concerning, however, and something that a semi-parametric approach would likely not resolve, is the disparity between the distributions of the most productivity years within the empirical data and in the simulation. The simulated data from this exponential model is much more likely to present its most productive year later in the career than the empirical data, and respectively underestimate the probability of an early-career peak, despite the fact that its inferred changepoint was so early in the career ($KS=0.30$, $p<0.001$; Fig. S8B).
XIV.2 Poisson-distributed maturation of synchronized projects
In a second simplified model, new faculty initiate $K$ projects at the same time, each of which finishes (yielding a paper) according to a Poisson distribution with rate $\lambda$. Once a project is finished, it resets. Simply through the synchronization and subsequent decoherence of the initial basket of projects, one can observe trajectories that appear almost canonical, for example, if faculty work on 10 projects each of which mature with rate 7 (Fig. S9A).
Figure S9: Second simplified model: Poisson-distributed maturation of synchronized projects. (A) Average productivity per year for simulated trajectories with $K=10$, $\lambda=7$. (B) Distribution of change in productivity increments compared with the fitted Laplace distribution for the (10, 7) model. (C) Average productivity per year for simulated trajectories with $K=5$, $\lambda=3$. (D) Distribution of change in productivity increments compared with the fitted Laplace distribution for the (5, 3) model.
However, in more sensible ranges of the parameter space, where a researcher embarks upon 5 projects and completes them on an average of 3 years, then the resulting trajectory is no longer canonical (Fig. S9C). Another discrepancy between this and empirical reality is that the productivity increments are far from symmetrically exponentially distributed (Fig. S9B,D). Despite the unsuitability of this second simplified model for modeling scientific productivity, we note it as an interesting and natural way that canonical trajectories may emerge in other contexts.
XV Relevance for prediction of individual trajectories
Predicting individual productivity would certainly serve as an impressive validation of any model, but our work outlines the challenge faced by any such predictor: simply by structuring the year-to-year variability of faculty productivity on top of a very basic random walk model, there is enough richness to reproduce the canonical trajectory, the diversity of individual trajectories, and the Shockley cumulative productivity distribution. In other words, many interesting phenomena about scientific productivity trajectories can be more parsimoniously explained by aspects of the variance, rather than the conditional expectations. Indeed, Ref. [48] performed a regression analysis of 12,400 Norwegian researchers’ productivity, controlling for gender, age, and position, and found that they could only explain 13.5 to 19% of the variance in productivity. Thus even with these crucial factors (and country, implicitly) controlled for, 80% of individual productivity remains unmodeled.
A focus on prediction collapses the variance into an average in two senses: first, the objective becomes to model the conditional expectation of productivity given covariates such as the prior productivity; second, the overall loss function, such as the root mean squared error, is typically an average of losses across individuals. Here our model is fundamentally more interested in predicting the overall distribution of productivity trajectories, including their outliers, rather than any individual trajectories alone. From the perspective of “predicting the population”, we already include many predictive model checks, since we generate simulated trajectories, and statistically compare those trajectories with the empirical ones. For instance, we find it interesting that our model successfully predicts the occurrence of outliers that were not explicitly included in the model, such as the distribution of faculty whose most productive year is very early or very late in their careers (Fig. 4B).
From the point of view of conditional expectations, our model expects regression to the mean within the structure of the truncated Laplace distribution, such that within DBLP, no researcher has an expected productivity (given their prior year’s productivity) greater than 23 papers per year. Yet within our dataset, over 772 researcher-years exhibit productivity greater than 23 papers, including 26 highly productive researcher-years with over 50 papers.
Our predictive model checks indicate that there is more inertia to faculty careers than our model would predict (Fig. 4A). For that reason, we may expect a simple baseline predictor where productivity is predicted to be exactly the same as the previous year to even outperform our model. Indeed, using 1000 bootstrap samples, we find that the average RMSE of the extrapolatory baseline across bootstraps is 837.96 (95% bootstrap CI: [827.19, 848.37]), while predictions from our model yield an average RMSE of 1054.09 (95% bootstrap CI: [1035.66, 1071.96]). These predictive checks remind us that our approach to capturing the overall distribution by only modeling global temporal structure currently comes at the expense of certain individual variance that can be retained.
When used for prediction, our model collapses the variance across individuals more than these simple baselines. If we apply a simple baseline of predicting a researcher’s output to be a constant average of their first few years of productivity (or their last observed productivity in the early career) then we have surrendered all within-individual variation, but preserved across-individual diversity. Our model, when deterministically applied using conditional expectations, performs a fixed point iteration (Fig. S4), thereby driving all faculty toward a constant productivity.